OpenStudy (anonymous):

(9x^4 + 3x^3y - 5x^2y ^2 + xy ^3) / (3x^3 + 2x^2y - xy^2)

OpenStudy (anonymous):

i can't understand you clearly

OpenStudy (anonymous):

5x^2y ^2 =\[5x^2y ^2 =5x^{2y ^{2}}\]

OpenStudy (anonymous):

or say \[5x^2y ^2 =5x^{2}y ^{2}\]

OpenStudy (anonymous):

oh. i see it's \[(9x ^{3}-y+y ^{3})/(3x ^{2}+4y-y ^{2})\]